Returns Year X Y 1 10 % 23 % 2 15 25 3 9 11 4...

Returns
Year X Y
1 10 % 23 %
2 15 25
3 9 11
4 14 15
5 10 15

  

Using the returns shown above, calculate the arithmetic average returns, the variances, and the standard deviations for X and Y. (Do not round intermediate calculations. Enter your average return and standard deviation as a percent rounded to 2 decimal places, e.g., 32.16, and round the variance to 5 decimal places, e.g., .16161.)
Answer #1

A B с D E F G Н. 1 х Y 2 Year 3 1 4 2 23% 25% 11% =(return of X-average return of X)^2 =(return of Y-average return of Y)^2 0

A B с D E F 1 2 Year Y 3 1 х 0.1 0.15 4 2 0.23 0.25 0.11 -0.15 =(return of X-average return of X)^2 =(return of Y-average ret